40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  CTN Lab elog  Not logged in ELOG logo
Entry  Tue Dec 16 21:31:21 2014, Kate, DailyProgress, , Minor changes to layout of cavity transmitted path 
    Reply  Wed Dec 17 20:30:36 2014, Aidan, Evan, DailyProgress, , Minor changes to layout of cavity transmitted path 
Message ID: 1536     Entry time: Tue Dec 16 21:31:21 2014     Reply to this: 1537
Author: Kate 
Type: DailyProgress 
Category:  
Subject: Minor changes to layout of cavity transmitted path 

Evan, Kate

We made a few minor modifications to the optical breadboard in transmission of the cavities.

For one, we removed the QWPs which were the first optics in the transmission paths. These had been necessary for the prior cavities where the Silica Tantala mirror coatings were not birefringent. The circular polarization which was transmitted needed to be turned into linear polarization to get the beat note on the PD. Now, because the cavities with AlGaAs coatings are birefringent, the resonant and transmitted light is already linearly polarized and the QWPs unnecessary. Before removing them, the power on the main readout PD, a PD1811, was 208 mV. Afterwards, it was 194 mV. 

Second, we started to set up the fiber coupler to send some light to the ATF lab where it will later be used in a PLL to stabilize the laser used for the seismometer sensing. There's a 29.5 uW pick-off of the North cavity transmitted light which had been dumped. I found a high reflector mirror to put in its place to direct light to the fiber. I also made sure the fiber coupler at the other end is secured to the table and the output dumped. A first attempt to couple the light did not work, but I need to find a way either to monito remotely the power transmitted or just temporarily feed the fiber back into the CTN lab. 

 

ELOG V3.1.3-